OpenStudy (bleuspectre):

Simplify square root of negative 72. negative 6 square root of 2 6 square root of negative 2 6 square root of 2 i 6 i square root of 2

OpenStudy (anonymous):

\[\sqrt{-72} = \sqrt{-1}\sqrt{72} \qquad i=\sqrt{-1} \]

OpenStudy (anonymous):

so what are the factors of 72?

OpenStudy (anonymous):

\[\sqrt{-72}\] This is the same as i\[\sqrt{72}\] 72 is not a perfect square root. i\[\sqrt{3*3*2*2*2}\]